OpenStudy (anonymous):

Use the Pythagorean Theorem to find the length of the missing side of the right triangle with both legs equal to √2. A. 2 B. 2 √2 C. √5 D. None of the above

OpenStudy (anonymous):



OpenStudy (anonymous):

For third side that is Hypotenuse too, you will use: \[c^2 = a^2 + b^2\] Here : \(c\) you want to find out. \(a = \sqrt{2}\) \(b = \sqrt{2}\) Just plug in the values.

OpenStudy (anonymous):

So is it B then?

OpenStudy (anonymous):

How?? No..

OpenStudy (anonymous):

LOL well idk I had D first then I B...

OpenStudy (anonymous):

Guessing is dangerous to your health.. Try avoiding it..

OpenStudy (anonymous):

Lol I know this...but I put in the values & whatevs and I didnt get it...

OpenStudy (anonymous):

Can you post your solution here so that I can check where you are getting wrong in this?

OpenStudy (anonymous):

\[(\sqrt{2})^2 = ??\]

OpenStudy (anonymous):

Are you trying to say it's A?

OpenStudy (anonymous):

Obviously..

OpenStudy (anonymous):

Lol why do you have to be so rude? I obvi didn't get it.